# V2Xum-LLM: Cross-Modal Video Summarization with Temporal Prompt Instruction Tuning

### v1

@@ Abstract
Video summarization aims to create short, accurate, and cohesive summaries of longer videos. Despite the existence of various video summarization datasets, a notable limitation is their limited amount of source videos, which hampers the effective fine-tuning of advanced large vision-language models (VLMs). Additionally, most existing datasets are created for video-to-video summarization, overlooking the contemporary need for multimodal video content summarization. Recent efforts have been made to expand from unimodal to multimodal video summarization, categorizing the task into three sub-tasks based on the summary’s modality: video-to-video (V2V), video-to-text (V2T), and a combination of video and text summarization (V2VT). However, the textual summaries in previous multimodal datasets are inadequate. To address these issues, we introduce Instruct-V2Xum, a cross-modal video summarization dataset featuring 30,000 diverse videos sourced from YouTube, with lengths ranging from 40 to 940 seconds and an average summarization ratio of 16.39%. Each video summary in Instruct-V2Xum is paired with a textual summary that references specific frame indexes, facilitating the generation of aligned video and textual summaries. In addition, we propose a new video summarization framework named V2Xum-LLM. V2Xum-LLM, specifically V2Xum-LLaMA in this study, is the first framework that unifies different video summarization tasks into one large language model’s (LLM) text decoder and achieves task-controllable video summarization with temporal prompts and task instructions. Experiments show that V2Xum-LLaMA outperforms strong baseline models on multiple video summarization tasks. Furthermore, we propose an enhanced evaluation metric for V2V and V2VT summarization tasks.
11footnotetext: Equal Contribution.22footnotetext: Corresponding author.
1 Introduction
@@
The interest in sharing life experiences has surged in recent years, making video the most informative and diverse visual medium on social media platforms. This trend has led to significant demands for a variety of video and language understanding tasks, such as video captioning [79], video question answering [82, 78], moment retrieval [32], and video summarization [14, 61]. Video summarization (V2V) provides an efficient way for humans to obtain key information from a long video. This process entails selecting the most significant information from a video and condensing it into a shorter form, all while maintaining the essence of the original content. Beyond extensive research in V2V summarization, there are a few recent explorations of V2T and V2VT summarization. Most notably, VideoXum [37] seeks to broaden the modality of video summaries to include text summaries. It utilizes the dense captions from ActivityNetCap [31] videos as text summaries and annotates the corresponding video segments as video summaries. Other datasets for video summarization include TVSum [61], SumMe [14], QFVS [57], MED Summaries [53], and so on. While it is expected that powerful LLMs can help improve video summarization, the insufficient number of source videos may not be able to support the robust fine-tuning of LLMs to perform this task since finetuning a large model with a limited number of training examples is prone to overfitting [20, 21]. In addition, the data compiled in VideoXum cannot be considered genuine summaries because the dense captions of ActivityNetCap contain a significant amount of redundant information. More recently, Shot2Story20K [15] collects 20k video-text data that enables the robust finetuning of LLMs, but it only supports video-to-text summarization. To address these issues, we propose Instruct-V2Xum, a new large-scale cross-model video summarization dataset that contains 30k open domain videos, partitioned as 25,000 in the training set, 1,000 in the validation set, and 4,000 in the test set. In Instruct-V2Xum, we obtain the source videos from YouTube using the video list provided by InternVid [74]. The methodology of video summarization parallels that of extractive text summarization, where the objective is to isolate the pivotal frames or sentences from the source videos or documents, respectively. Extractive text summarization is a foundational task in the NLP field, with numerous LLMs setting the benchmark in this domain [85, 26, 17]. Inspired by this, we first extract frames from the source videos and employ LLaVA-1.5-7B [39] to generate detailed captions for each frame. Then we take all the frame captions as a document to perform extractive document summarization using GPT-4 [1]. This approach enables us to obtain both video summaries and their corresponding textual summaries. Finally, the extracted text summaries are further refined into a grammar-fluent short paragraph by GPT-4.
@@
Numerous visual instruction tuning-based methods have been proposed for video-language understanding [22, 70, 36]. These models can process the long videos for general video-language understanding and reasoning tasks such as video question answering, video captioning, and so on. Recently, there have been some attempts for fine-grained video moments understanding [22] or video-language temporal grounding [38] using large VLMs. However, these approaches, which typically process video frames as sequential images for a frozen visual encoder and train an LLM decoder for identifying video moment boundaries, are not well-suited for dense temporal prediction in video summarization tasks, particularly in the V2VT tasks. Furthermore, most existing models require large-scale data to train new parameters added to pre-trained VLMs [38, 36, 44]. This requirement significantly limits their practicality in scenarios where only a limited number of training examples are available. To address these problems, we design a new temporal prompt instruction tuning framework – V2Xum-LLaMA. In V2Xum-LLaMA, we unify different modalities of video summary generation into one LLM decoder. This framework stands out by removing the dependence on task-specific layers that were required for video summarization in earlier VLM-based approaches. The main advantages of this method are that it enables the effective adaptation of the learned knowledge and the powerful capabilities of the pretrained language models to dense video temporal and content understanding. All the pretrained parameters of the VLMs are reused, and the model takes interleaved video frames and natural language temporal prompts as input to facilitate end-to-end model training. As video temporal prediction is performed by using language models, there is a challenge for calculating the correlation for V2V evaluation, since the language model-predicted video summaries are the discrete frame indexes. To overcome this challenge, we propose a solution in section 5 to calculate the scores for language model-predicted video summaries. Furthermore, we also provide an analysis of the existing video summarization tasks and propose $F_{CLIP}$ and $Cross-F_{CLIP}$, the enhanced evaluation metrics for V2V and V2VT summarization tasks.
@@
  • We propose a novel and flexible cross-modal video summarization framework – V2Xum-LLaMA, where we unify different video summarization tasks into one pre-trained language decoder without introducing the randomly initialized task-specific heads. The proposed framework takes interleaved video frames and natural language temporal prompts as input, which facilitates the pre-trained VLMs to process long sequences in an end-to-end manner. Our proposed method outperforms all the selected strong baseline models on existing mainstream V2V, V2T, and V2VT benchmarks.
@@
  • To address the challenge of insufficient video-language data for effectively fine-tuning large VLMs for various video summarization tasks, we build a new instruction following dataset named Instruct-V2Xum for cross-model video summarization. This dataset comprises 30k diverse videos sourced from YouTube, with durations spanning from 40 to 940 seconds. The proposed dataset enables VLMs to generate modality-controllable video summaries with the task prompt. The experiments validate the rationality of our proposed dataset.
@@
  • We provide a comprehensive analysis of various existing video summarization tasks from the perspectives of data, method, and evaluation. Building on this analysis, we propose $F_{CLIP}$ and $Cross-F_{CLIP}$, an enhanced evaluation metric for V2V and V2VT summarization tasks. Experimental results indicate that our proposed evaluation metrics are highly consistent with the traditional evaluation metrics including F1, Spearman correlation, and Kendall correlation.
2 Related Work
2.1 Video Summarization
Video summarization is the process of distilling salient information from video content by analyzing its structure and temporal redundancies. Traditional video summarization, also known as video-to-video summarization, typically generates a condensed version of the original video, comprising selected frames [41, 13, 71, 8], shots [25, 9, 88], or segments [66, 30]. These models are commonly trained using supervised learning approaches, with reinforcement learning methods like policy gradient [77], optimizing for diversity and representativeness in the summarized output [91], gaining popularity. The standard datasets for video summarization tasks include SumMe [14] and TVSum [61], which are widely used for benchmarking purposes. In recent years, cross-modal video summarization [11, 16, 24] has emerged as an area of interest, incorporating additional modalities such as audio, speech, subtitles, and captions. This approach leverages multimodal models to create more comprehensive summaries. Video-to-text summarization [50, 4] is an evolving field that aims to generate descriptive paragraphs in natural language that encapsulate video content. VideoXum [37] makes progress in this direction by repurposing the ActivityNetCap [31] dataset and employing the vision-language model BLIP-2 [33] for both video-to-video and video-to-text summarization. However, it is important to note that the data collected in [37] may not constitute genuine summaries, as the dense captions from the ActivityNetCap dataset often contain significant redundant information.
2.2 Video Temporal Understanding
Beyond video summarization, there has been an increasing interest in temporal understanding tasks for video analysis. Temporal localization plays a pivotal role in comprehending long-form or untrimmed videos by associating specific segments with their corresponding semantics. Research has primarily concentrated on tasks such as temporal video grounding [42, 75], dense video captioning [72, 81], video highlight detection [32, 28], step localization [58], and procedure planning [2]. Despite significant progress, models designed for action temporal localization and highlight detection often rely on predefined labels [83], which hampers a comprehensive understanding. To address these limitations, research has focused on generic event boundary detection [59, 73, 67] and dense video captioning [31], utilizing datasets that provide event descriptions [37]. Nevertheless, Most existing models still rely on regression for temporal predictions, necessitating separate heads for captioning and regression tasks [72, 86, 66]. Recent advancements in Large Language Models (LLMs) have introduced a paradigm shift by enabling the direct use of natural language to specify temporal locations in videos, thus offering a more intuitive approach to video understanding.
2.3 Large Language Models
In recent years, Large Language Models (LLMs) have witnessed rapid advancements [1, 68, 69, 89]. With pretraining on extensive corpora from the Internet, LLMs acquire substantial knowledge, enabling powerful zero-shot and in-context learning capabilities [1, 76, 29, 19]. Efforts have been increasingly directed toward leveraging LLMs for multimodal tasks [43, 60, 63]. Techniques such as vision-language alignment and adapter fine-tuning are employed to integrate LLMs into the multimodal domain. These methods align the visual features extracted by visual encoders with the input token space of LLMs [39]. Typically, the parameters of LLMs are frozen to retain their existing capabilities although LoRA [18] fine-tuning is occasionally applied to adapt to data with varying distributions or to acquire new behaviors. Utilizing this approach, several studies [70, 44, 36, 84] have successfully employed LLMs for video understanding tasks, referred to as Vid-LLMs [64]. However, current research primarily concentrates on general video understanding tasks, such as video question-answering (QA) and video captioning, with less emphasis on temporal information. Recent works [22, 55, 65] have started exploring LLMs’ potential in temporal grounding and localization, highlighting the untapped potential of LLMs in video temporal understanding.
Table 1: Comparison with existing video-to-video summarization and video-to-text summarization datasets. “V2V”, “V2T”, and “V2VT” denote whether the datasets support video-to-video summarization, video-to-text summarization, or both tasks, respectively. “Instruction” denotes whether the dataset supports video-text instruction tuning. M and S stand for manual and model synthesized, respectively.
@@ 2.3 Large Language Models
MSVD [3] Open 1,970 M
YouCook [5] Cooking 88 M
MSR-VTT [79] Open 7,180 M
UCF101 [62] Open 13,320 M
ActivityNetCap [31] Activities 20,000 M
Shot2Story20k [15] Open 20,000 M+S
SumMe [14] Events, holidays, sports 25 M
TVSum [61] News, documentaries, vlogs 50 M
OVP [6] Documentaries, lectures 50 M
VSUMM [7] Cartoons, news, commercials, shows 50 M
EDUVSUM [12] Letures 98 M
LoL [10] Matches of League of Legends 218 M
Ads-1K [66] Commercials 1,041 M+S
VideoXum [37] Activities 14,001 M
@@
3 The Instruct-V2Xum Dataset
In this section, we introduce the details of data collection and provide the quantity and quality analysis for Instruct-V2Xum, which is an instruction-tuning dataset for cross-modal Video summarization.
3.1 Data Curation
3.1.1 Frame Captioning and Extractive Summarization.
@@ 3.1 Data Curation
The source videos are sampled from InternVid [74]. We first filter the raw videos according to the duration of the videos and their aesthetic scores. We obtain 30k for video and text summaries generation. Initially, we extract frames at a rate of 1 FPS, followed by the transformation of these sampled video frames into their detailed textual representations using LLaVA-1.5-7B [39]. Upon obtaining the text representations of the video frames, we utilize GPT-4 to perform extractive document summarization. Finally, the frame captions from the text summaries are converted into both video and text summaries.
3.1.2 Text Summarization Refinement.
@@
To further reduce the redundancy of the text summaries, we utilize BERT score [90] to filter out the frame text representations that are similar to other summary frames. Here, we set the threshold to 0.93. The filtered video frame captions’ indexes serve as the video summaries for the source videos. Then, we employ GPT-4 to further compress and rewrite the video summaries to be shorter and more grammar-fluent.
3.2 Quantity Analysis
@@
Here we provide the statistical results for our collected data. Figure 2 illustrates the distribution of the duration of source videos and the summaries length for both video and text in Instruct-V2Xum. We also provide more statistical results in the appendix. The average duration of the source videos is 183 seconds. The average length of the text summaries is 239 tokens. The average length of the video summaries is 30 frames. The average compression ratio is 16.39%.
3.3 Quality Analysis
@@
Since the text summaries are synthesized by LLMs, we analyze the quality of the generated summaries from the perspective of grammar fluency and commonsense plausibility. For the grammar-fluency evaluation, we leverage a grammar-check model [46] that assigns high scores to correct texts grammatically. For the commonsense plausibility evaluation, we utilize Vera [40], a plausibility estimation model, to identify the nonsensical bias. The results of these evaluations are presented in the appendix. Overall, our data demonstrates reasonable scores in both grammar and Vera evaluations.
Figure 3: The architecture of our proposed V2Xum-LLaMA.
4 Methodology
4.1 Overview
This section presents our proposed unified cross-modal video summarization framework V2Xum-LLaMA. Our method is distinguished by its utilization of temporal prompt encoding and temporal-aware decoding to condense and summarize video content effectively. As depicted in Fig 3, our model architecture comprises three key components: video frames are encoded through a vision encoder, subsequently combined with temporal prompts, and finally processed by a decoder-only LLM. Utilizing LLaMA as the decoder, our system is capable of generating both video-to-video and video-to-text summarizations. Temporal Prompt Encoding associates video frames with corresponding timestamps, embedding positional information into each frame. These encodings are then merged with specific summarization instructions, and the summaries are produced through temporal-aware decoding via LLaMA. Our methodology allows for modality control over the generated summaries, based on the provided instructions, producing either video or text summarizations, or a combination thereof. This innovative encoding-decoding mechanism enhances the model’s understanding of the video timeline and improves the summarization output’s relevance and accuracy.
4.2 Interleaved Video and Temporal Prompt Encoding
@@
The temporal prompt mechanism is designed to bind the visual tokens with the corresponding timestamps at the frame level to inject positional information for each frame. To this end, we first encode each video frame $f_{i}$ using pretrained CLIP [54] encoder $E_{v}$. The frames’ encoding $v_{i}=E_{v}(f_{i})$ are involved in the visual tokens sequence $V=\{v_{1},v_{2},...,v_{L}\}$, where $L$ is the number of sampled frames.
@@
We then bind temporal prompts with each visual token. The temporal prompts are tokenized zero-padded numbers in natural language format like “00”, “06”, “12”, “99”, etc., indicated by $T=\{t_{1},t_{2},...,t_{L}\}$. They are inserted into the visual token sequence $V$ to form a new sequence with interleaved visual tokens and temporal prompts:
$S=\{t_{1},v_{1},t_{2},v_{2},...,t_{L},v_{L}\}.$
@@
Compared to the original visual token sequence, the temporal prompted sequence can better capture the relations between the timestamps and the visual semantics.
4.3 Temporal-Aware Decoding
@@
We use decoder-only LLM, LLaMA [69] as the decoder to generate video-to-video summarization $A^{v}$, video-to-text summarization $A^{t}$, or both $A^{b}$. The type of generated summarization is controlled by the instructions $I^{t}$, $I^{v}$, or $I^{b}$ that request the model to generate video summarization, text summarization, or both, respectively. With the instruction $I$ and temporal prompted sequence $S$, the output is given by the temporal-aware decoding:
$A^{x}=LLM(S,I^{x}),~{}x\in\{v,t,b\}.$
@@
The video-to-video summarization $A^{v}$ is defined as a sequence of temporal tokens sequence, i.e., $A^{t}=\{t_{i}\}^{M}_{i=1}$, which is a subset of temporal prompts sequence $T$ and $M\leq L$. The video-to-text summarization is a summarized caption in a natural language format. In our implementation shown as Fig. 3, the video-to-text summarization also contains frame referring temporal tokens, i.e.:
$A^{t}=\{...,w_{i-1},t_{j},w_{i+1},w_{i+2}...\},$

### v2

@@ Abstract
Video summarization aims to create short, accurate, and cohesive summaries of longer videos. Despite the existence of various video summarization datasets, a notable limitation is their limited amount of source videos, which hampers the effective training of advanced large vision-language models (VLMs). Additionally, most existing datasets are created for video-to-video summarization, overlooking the contemporary need for multimodal video content summarization. Recent efforts have been made to expand from unimodal to multimodal video summarization, categorizing the task into three sub-tasks based on the summary’s modality: video-to-video (V2V), video-to-text (V2T), and a combination of video and text summarization (V2VT). However, the textual summaries in previous multimodal datasets are inadequate. To address these issues, we introduce Instruct-V2Xum, a cross-modal video summarization dataset featuring 30,000 diverse videos sourced from YouTube, with lengths ranging from 40 to 940 seconds and an average summarization ratio of 16.39%. Each video summary in Instruct-V2Xum is paired with a textual summary that references specific frame indexes, facilitating the generation of aligned video and textual summaries. In addition, we propose a new video summarization framework named V2Xum-LLM. V2Xum-LLM, specifically V2Xum-LLaMA in this study, is the first framework that unifies different video summarization tasks into one large language model’s (LLM) text decoder and achieves task-controllable video summarization with temporal prompts and task instructions. Experiments show that V2Xum-LLaMA outperforms strong baseline models on multiple video summarization tasks. Furthermore, we propose an enhanced evaluation metric for V2V and V2VT summarization tasks.
Introduction
@@
The interest in sharing life experiences has surged in recent years, making video the most informative and diverse visual medium on social media platforms. This trend has led to significant demands for a variety of video and language understanding tasks, such as video captioning (Xu et al. 2016), video question answering (Yu et al. 2019; Xiao et al. 2021), moment retrieval (Lei, Berg, and Bansal 2021), and video summarization (Gygli et al. 2014; Song et al. 2015). Video summarization (V2V) provides an efficient way for humans to obtain key information from a long video. This process entails selecting the most significant information from a video and condensing it into a shorter form, all while maintaining the essence of the original content. Beyond extensive research in V2V summarization, there are a few recent explorations of V2T and V2VT summarization. Most notably, VideoXum (Lin et al. 2023b) seeks to broaden the modality of video summaries to include text summaries. It utilizes the dense captions from ActivityNetCap (Krishna et al. 2017) videos as text summaries and annotates the corresponding video segments as summaries. Other datasets for video summarization include TVSum (Song et al. 2015), SumMe (Gygli et al. 2014), QFVS (Sharghi, Laurel, and Gong 2017), MED Summaries (Potapov et al. 2014), and so on. While it is expected that powerful LLMs can help improve video summarization, the insufficient number of source videos may not be able to support the robust fine-tuning of LLMs to perform this task since finetuning a large model with a limited number of training examples is prone to overfitting (Hua et al. 2021, 2023). In addition, VideoXum data cannot be considered true summaries due to redundant information in ActivityNetCap’s dense captions. More recently, Shot2Story20K (Han et al. 2023) collects 20k video-text data that enables the robust finetuning of LLMs, but it only supports video-to-text summarization. To address these issues, we propose Instruct-V2Xum, a new large-scale cross-model video summarization dataset that contains 30k open domain videos, partitioned as 25,000 in the training set, 1,000 in the validation set, and 4,000 in the test set. In Instruct-V2Xum, we obtain the source videos from YouTube using the video list provided by InternVid (Wang et al. 2023b). The methodology of video summarization parallels that of extractive text summarization, where the objective is to isolate the pivotal frames or sentences from the source videos or documents, respectively. Extractive text summarization is a foundational task in the NLP field, with numerous LLMs setting the benchmark in this domain (Zhang, Liu, and Zhang 2023; Jia et al. 2020; He et al. 2023a). Inspired by this, we first extract frames from the source videos and employ LLaVA-1.5-7B (Liu et al. 2023a) to generate detailed captions for each frame. Then we take all the frame captions as a document to perform extractive document summarization using GPT-4 (Achiam et al. 2023). This approach enables us to obtain both video summaries and their corresponding textual summaries. Finally, the extracted text summaries are further refined by GPT-4.
@@
Numerous visual instruction tuning-based methods have been proposed for video-language understanding (Huang et al. 2023; Wang et al. 2023a; Lin et al. 2023a). These models can process the long videos for general video-language understanding and reasoning tasks such as video question answering, video captioning, and so on. Recently, there have been some attempts for fine-grained video moments understanding (Huang et al. 2023) or video-language temporal grounding (Lin et al. 2023c) using large VLMs. However, these approaches, which typically process video frames as sequential images for a frozen visual encoder and train an LLM decoder for identifying video moment boundaries, are not well-suited for dense temporal prediction in video summarization tasks, particularly in the V2VT tasks. Furthermore, most existing models require large-scale data to train new parameters added to pre-trained VLMs (Lin et al. 2023c, a; Maaz et al. 2023; He et al. 2023b). This requirement significantly limits their practicality in scenarios where only a limited number of training examples are available. To address these problems, we design a new temporal prompt instruction tuning framework – V2Xum-LLaMA. In V2Xum-LLaMA, we unify different modalities of video summary generation into one LLM decoder. This framework stands out by removing the dependence on task-specific layers that were required for video summarization in earlier VLM-based approaches. The main advantages of this method are that it enables the effective adaptation of the learned knowledge and the powerful capabilities of the pretrained language models to dense video temporal and content understanding. All the pretrained parameters of the VLMs are reused, and the model takes interleaved video frames and natural language temporal prompts as input to facilitate end-to-end model training. As video temporal prediction is performed by using language models, there is a challenge for calculating the correlation for V2V evaluation, since the language model-predicted video summaries are the discrete frame indexes. To overcome this challenge, we propose a solution to calculate the scores for language model-predicted video summaries. Furthermore, we also provide an analysis of the existing video summarization tasks and propose $F_{CLIP}$ and $Cross-F_{CLIP}$, the enhanced evaluation metrics for V2V and V2VT summarization tasks.
@@
  • We propose V2Xum-LLaMA, a novel cross-modal video summarization framework that unifies different tasks into a single pre-trained language decoder, eliminating the need for task-specific heads used in prior methods. By taking interleaved video frames and temporal prompts as input, our method enables end-to-end processing of long video sequences and outperforms all strong baseline models on mainstream V2V, V2T, and V2VT benchmarks.
@@
  • To address the lack of video-language data for fine-tuning large VLMs in video summarization tasks, we created Instruct-V2Xum, a new instruction-following dataset for cross-modal video summarization. It contains 30k diverse YouTube videos, ranging from 40 to 940 seconds, enabling VLMs to generate modality-controllable video summaries with task prompts. The experiments validate the rationality of our proposed dataset.
@@
  • We present a comprehensive analysis of the limitations in current video summarization tasks from the perspectives of data, methods, and evaluation. Based on this, we propose $F_{CLIP}$ and $Cross-F_{CLIP}$, an enhanced evaluation metric for V2V and V2VT summarization tasks. Experimental results show that these metrics are highly consistent with the traditional evaluation metrics including F1, Spearman correlation, and Kendall correlation.
Related Work
@@
MSVD (Chen and Dolan 2011) Open 1,970 M
YouCook (Das et al. 2013) Cooking 88 M
MSR-VTT (Xu et al. 2016) Open 7,180 M
UCF101 (Soomro, Zamir, and Shah 2012) Open 13,320 M
ActivityNetCap (Krishna et al. 2017) Activities 20,000 M
Shot2Story20k (Han et al. 2023) Open 20,000 M+S
SumMe (Gygli et al. 2014) Events, holidays, sports 25 M
TVSum (Song et al. 2015) News, documentaries, vlogs 50 M
OVP (de Avila et al. 2011) Documentaries, lectures 50 M
VSUMM (De Avila et al. 2011) Cartoons, news, commercials, shows 50 M
EDUVSUM (Ghauri, Hakimov, and Ewerth 2020) Letures 98 M
LoL (Fu, Tai, and Chen 2019) Matches of League of Legends 218 M
Ads-1K (Tang et al. 2022) Commercials 1,041 M+S
VideoXum (Lin et al. 2023b) Activities 14,001 M
@@
Table 1: Comparison with existing video-to-video summarization and video-to-text summarization datasets. “V2V”, “V2T”, and “V2VT” indicate support for video-to-video, video-to-text, or both tasks. “Instruction” denotes whether the dataset supports video-text instruction tuning. M and S stand for manual and model synthesized, respectively.
Video Summarization
Traditional video summarization, also known as video-to-video summarization, typically generates a condensed version of the original video, comprising selected frames (Liu, Li, and Wang 2020; Ghauri, Hakimov, and Ewerth 2021; Wang et al. 2019; Fajtl et al. 2019), shots (Ji et al. 2019; Feng et al. 2018; Zhang, Grauman, and Sha 2018), or segments (Tang et al. 2022; Koutras and Maragos 2019). These models are commonly trained using supervised learning approaches, with reinforcement learning methods like policy gradient (Williams 1992), optimizing for diversity and representativeness in the summarized output (Zhou, Qiao, and Xiang 2018), gaining popularity. The standard datasets for video summarization tasks include SumMe (Gygli et al. 2014) and TVSum (Song et al. 2015), which are widely used for benchmarking purposes. In recent years, cross-modal video summarization (Fu, Wang, and Yang 2020; Haopeng et al. 2022; Huang et al. 2021) has emerged as an area of interest, incorporating additional modalities such as audio, speech, subtitles, and captions. These approaches leverage multimodal models to create more comprehensive summaries. Video-to-text summarization (Palaskar et al. 2019; Choi, Oh, and Kweon 2018) is an evolving field that aims to generate descriptive paragraphs in natural language that encapsulate video content. VideoXum (Lin et al. 2023b) advances this field by repurposing the ActivityNetCap (Krishna et al. 2017) dataset and employing the BLIP-2 (Li et al. 2023) model for both V2V and V2T summarization. However, the data in (Lin et al. 2023b) may not be genuine summaries, as the dense captions from ActivityNetCap often include significant redundancy.
Large Language Models
In recent years, Large Language Models (LLMs) have witnessed rapid advancements (Achiam et al. 2023; Touvron et al. 2023a, b; Zhang et al. 2023). With pretraining on extensive corpora from the Internet, LLMs acquire substantial knowledge, enabling powerful zero-shot and in-context learning capabilities (Achiam et al. 2023; Wei et al. 2022; Kaplan et al. 2020; Hu et al. 2022b). Efforts have been increasingly directed toward leveraging LLMs for multimodal tasks (Lyu et al. 2023; Shu et al. 2023; Su et al. 2023; Yu et al. 2024; Hua et al. 2024). Techniques such as vision-language alignment and adapter fine-tuning are employed to integrate LLMs into the multimodal domain. These methods align the visual features extracted by visual encoders with the input token space of LLMs (Liu et al. 2023a). Typically, the parameters of LLMs are frozen to retain their existing capabilities although LoRA (Hu et al. 2022a) fine-tuning is sometimes applied in low-resource settings. Based on this, several studies (Wang et al. 2023a; Maaz et al. 2023; Lin et al. 2023a; Zhang, Li, and Bing 2023) have successfully employed LLMs for video understanding tasks, referred to as Vid-LLMs (Tang et al. 2023). However, current research primarily concentrates on general video understanding tasks, such as video question-answering (QA) and video captioning, with less emphasis on temporal information. Recent works (Huang et al. 2023; Ren et al. 2023; Tang et al. 2024) explore LLMs’ potential in temporal grounding and localization, emphasizing their untapped capability in video temporal understanding.
The Instruct-V2Xum Dataset
Data Curation
Frame Captioning and Extractive Summarization.
@@
The source videos are sampled from InternVid (Wang et al. 2023b). We first filter the raw videos according to their duration and the aesthetic scores. The filtered data is then used to generate both video and text summaries. Specifically, we extract video frames at a rate of 1 FPS and then convert these frames into detailed textual descriptions using LLaVA-1.5-7B. After obtaining the textualized video frames, we utilize GPT-4V to perform extractive document summarization. Finally, the extracted summaries are converted into coherent video and text summaries.
Figure 2: The architecture of the proposed V2Xum-LLaMA.
Text Summarization Refinement.
@@
To further reduce the redundancy of the text summaries, we utilize BERT score (Zhang et al. 2019) to filter out the frame text representations that are similar to other summary frames. Here, we set the threshold to 0.93. The filtered video frame captions’ indexes serve as the video summaries for the source videos. Then, we employ GPT-4 to further compress and rewrite the video summaries to be shorter and more grammar-fluent.
Human Verification.
To enhance the quality of the collected data, we employed human annotators to filter the GPT-4-generated data, resulting in a final set of 30k data points. We also show examples of human-filtered data in the Appendix.
Quantity Analysis
@@
We provide the statistical results for our collected data. Figure 6 in the appendix illustrates the distribution of the duration of source videos and the summaries length for both video and text in Instruct-V2Xum. We also provide more statistical results in the appendix. The average duration of the source videos is 183 seconds. The average text summary length is 239 tokens, and the average video summary length is 30 frames. The average compression ratio is 16.39%.
Quality Analysis
@@
We analyze the quality of the text summaries from the perspective of grammar fluency and commonsense plausibility. For the grammar-fluency evaluation, we leverage a grammar-check model (Morris et al. 2020) that assigns high scores to correct texts grammatically. For the commonsense plausibility evaluation, we utilize Vera (Liu et al. 2023b), a plausibility estimation model, to identify the nonsensical bias. The results of these evaluations are presented in the appendix. In summary, our human-filtered data demonstrates reasonable scores in both grammar and Vera evaluations.
Methodology
Overview
This section introduces our unified cross-modal video summarization framework, V2Xum-LLaMA, which employs the interleaved video frames along with temporal and task prompts as input and converts videos into multimodal summaries. Figure 2 depicts the architecture of our framework, where video frames are first encoded into visual representations, then combined with temporal and task prompts to obtain the interleaved input tokens for the LLaMA-2 decoder. Using the LLaMA decoder, our system can generate both video-to-video and video-to-text summaries. In addition, we design a temporal-aware decoding training strategy to enhance the alignment of text and video summaries. To sum up, our approach enables modality control, producing either video, text summaries, or both, and enhances the model’s understanding of the video timeline for more relevant and accurate outputs.
Interleaved Video and Temporal Prompt Encoding
@@
The temporal prompt mechanism binds visual tokens to their corresponding frame-level timestamps, injecting positional information for each frame. To this end, we first encode each video frame $f_{i}$ using pretrained CLIP (Radford et al. 2021) encoder $E_{v}$. The frames’ encoding $v_{i}=E_{v}(f_{i})$ are involved in the visual tokens sequence $V=\{v_{1},v_{2},...,v_{L}\}$, where $L$ is the number of sampled frames.
@@
We then bind temporal prompts with each visual token. The temporal prompts are tokenized zero-padded numbers in natural language format like “[f00]”, “[f06]”, “[f12]”, “[f99]”, etc., indicated by $T=\{t_{1},t_{2},...,t_{L}\}$. They are inserted into the visual token sequence $V$ to form a new sequence with interleaved visual tokens and temporal prompts:
$S=\{t_{1},v_{1},t_{2},v_{2},...,t_{L},v_{L}\}$ (1)
@@
Compared to the original visual token sequence, the temporal prompted sequence can better capture the relations between the timestamps and the visual semantics. Then the sequence $S$ is projected into the word embedding space by the vision adapter, which is represented as $\overline{S}$.
Temporal-Aware Decoding
@@
We use LLaMA-2 (Touvron et al. 2023b) as the decoder to generate V2V summarization $A^{v}$, V2T summarization $A^{t}$, or V2VT summarization $A^{b}$. The $I^{v}$, $I^{t}$, and $I^{b}$ represent the task instructions for V2V, V2T, and V2VT summarization. With the instruction $I$ and temporal prompted sequence $\overline{S}$, the output is given by the temporal-aware decoding:
$A^{x}=LLM(\overline{S},I^{x}),~{}x\in\{v,t,b\}$ (2)
@@
The V2V summarization $A^{v}$ is defined as a sequence of temporal tokens sequence, i.e., $A^{v}=\{v_{i}\}^{M}_{i=1}$, which is a subset of temporal prompts sequence $T$ and $M\leq L$. The V2T summarization is a summarized caption in a natural language format. In our implementation shown as Figure 2, the V2T summarization also contains frame referring temporal tokens:
$A^{t}=\{...,w_{i-1},t_{j},w_{i+1},w_{i+2}...\}$ (3)